A car accelerates from 4.0 m/s to 24 m/s at a rate of 2.5m2. How far does it travel while accelerating?

A) 69m B) 207m C) 41m D) 112m E) 980m

Equations to chose from:

x=xo+Vof +1/2at2

V=vo+at

V2=Vo2+2a(x-xo)
Answer #1

Using equation,

V^2 = u^2 + 2a(x - xo)

24^2 = 4^2 + 2 x 2.5 x X

24^2 - 4^2 = 5 x

X = 112 m

Comment in case any doubt please rate my answer...